# Universal Speech Token Learning via Low-Bitrate Neural Codec and Pretrained Representations

### v1

@@ Abstract
Index Terms:
Speech language model, neural speech codec, speech synthesis.
publicationid: pubid: $0$–$00$ 2024 IEEE
@@ III-B1 Encoder
Instead of training from scratch, we leverage a self-supervised learned semantic encoder for $E_{s}$ and extract residue features through another $E_{p}$ conditioned on intermediate features of $E_{s}$ and $\bm{S}$ (see Fig. 2). Specifically, we initialize $E_{s}$ from a pretrained multilingual HuBERT [5], i.e. mHuBERT Base  [^1] https://dl.fbaipublicfiles.com/hubert/mhubert_base_vp_en_es_fr_it3.pt trained on a 100K subset of the large multilingual corpus VoxPopuli. As the bottom or middle layers of HuBERT capture more acoustic information of speech and the top layers tend to capture more semantic information, we take the $11$th layer output of mHuBERT as our semantic feature $\bm{S}_{se}\ {\in\mathbb{R}^{T\times C_{a}}}$ as that in [19] and the $6$th layer of mHuBERT as our acoustic feature $\bm{A}_{ac}\ {\in\mathbb{R}^{T\times C_{a}}}$ to learn the local pair  $(\bm{S},\bm{P})$, $\bm{S}\in\mathbb{R}^{T\times C_{s}},\bm{P}\in\mathbb{R}^{T\times C_{p}}$.  Here $C_{a}$, $C_{s}$ and $C_{p}$ are channel dimensions and $T$ is the number of $40$ms frames. As our quantization is performed on a $40$ms frame to produce a more compact token sequence while mHuBERT features are in a $20$ms basis, a reshape to $40$ms by stacking two adjacent frames is conducted before extracting $\bm{S}_{se}$ and $\bm{A}_{ac}$. We should emphasize that as our focus is not on how to learn semantic features, but instead on investigating universal tokens capturing various attributes of speech with compactness and disentanglement that can benefit speech language modeling, here we just use existing mHuBERT in semantic encoder.
@@ III-B3 Decoder
  • An acoustic decoder $D_{a}$, which converts the generated acoustic tokens $\hat{\bm{A}}$ back into the waveform, as $\hat{x}=D_{a}(\hat{\bm{A}})$.
@@
Acoustic Tokens with TF-Codec We use the pretrained TF-Codec [20], a state-of-the-art low-latency neural speech codec, to extract the acoustic token sequence $\bm{A}$. TF-Codec follows a convolutional encoder-VQ-decoder paradigm, operating in the time-frequency domain. For simplicity, we disable the predictive loop in TF-Codec and use the non-predictive pretrained TF-Codec at 6 kbps that achieves high-fidelity speech reconstruction. Unlike RVQ, it uses group vector quantization  (GVQ) where the input features stacked at 50 Hz are split into $G_{a}=16$ groups along the channel dimension and each group is quantized by a separate vector quantizer with $1024$ codewords,  yielding the acoustic token sequence $\bm{A}=[\bm{a}_{0},\bm{a}_{1},...,\bm{a}_{T_{a}-1}]^{\mathsf{T}}\in\mathbb{R}^{T_{a}\times G_{a}}$. $\bm{a}_{t}=[a_{t}^{0},a_{t}^{1},...,a_{t}^{G_{a}-1}]^{\mathsf{T}}$ denotes the acoustic codes at $t$-th frame with $G_{a}$ groups and each element in $\bm{a}_{t}$ belongs to the set $\{0,1,2,...,1023\}$. $T_{a}$ is the number of acoustic frames.
@@
where $\bm{a}_{-1}$ is the $start$ token and $\bm{a}_{T_{a}}$ is the $eos$ token.  $\bm{a}_{<t}$ denotes all past frames from $\bm{a}_{-1}$ to $\bm{a}_{t-1}$ and we follow this notation in all equations below if not otherwise specified. To achieve this, the $G_{a}$ embeddings corresponding to $G_{a}$ codebooks are concatenated  along the channel dimension as the transformer input and $G_{a}$ classification heads are employed to predict the $G_{a}$ acoustic codes  in $\bm{a}_{t}$ jointly by projecting the transformer’s final hidden state to $G_{a}$ sets of logits, one for each of the $G_{a}$ codebooks.  It is noted that as the number of frames for $\bm{A}$ and $F_{fused}$ are different, i.e. $T_{a}=2T$, a reshape of $F_{fused}$ from $\mathbb{R}^{T\times{C_{a}}}$ to $\mathbb{R}^{T_{a}\times\frac{C_{a}}{2}}$ is conducted before being taken as the condition in the transform decoder.
@@ IV-A Zero-shot Text-to-Speech Synthesis (TTS)
$(\hat{\bm{S}},\hat{\bm{P}})=UniCodecLM(text;\tilde{x}),\hat{\bm{G}}=E_{g}(\tilde{x}),$ (11)
where $\tilde{x}$ is a speech prompt and $text$ is the text input. The $UniCodecLM$ is an autoregressive codec language model conditioned on $text$ and $\tilde{x}$, formulated as
@@
where $\bm{c}_{t}=[\bm{s}_{t},\bm{p}_{t}]$ denotes the concatenated semantic tokens $\bm{s}_{t}$ and residual tokens $\bm{p}_{t}$ of the $t$-th frame in $\bm{S}$ and $\bm{P}$ along the channel dimension.  $N$ denotes the length of UniCodec token sequence, and $\bm{c}_{-1}$ and $\bm{c}_{N}$ denote the $start$ and $eos$ tokens, respectively. $\bm{Z}_{phoneme}$ is the phoneme transcription of the input text. We use a decoder-only transformer with 12 layers, 16 heads and a dimension of 768 for $\theta_{TTS}$. For better speech intelligibility of the synthesized speech, we assign higher weights to the loss associated with $\bm{s}_{t}$ to direct the model’s focus more on predicting semantic information. With predicted triplet token from Eq. (11), the final speech is synthesized by the UniCodec decoder give by $\hat{x}=UniCodecDec(\hat{\bm{G}},\hat{\bm{S}},\hat{\bm{P}})$.
@@ IV-B Speech-to-Speech Translation (S2ST)
$\mathcal{L}_{S2ST}=-\ {\sum\limits_{t=0}^{N}\log}p(\bm{c}_{tgt,t}|\bm{c}_{tgt,{<t}},\bm{C}_{src},\bm{L}_{src},\bm{L}_{tgt};\theta_{S2ST}),$ (14)
@@ V-A Speech Resynthesis
In UniCodec, we set the target bitrate of $\bm{S}$ and $\bm{P}$ tokens to be 250 bps, respectively, leading to a total of 500 bps.  The group number $M$, the number of codewords $K$ in GVQ are set to 2 and 256, respectively, with both $R^{s}_{target}$ and $R^{p}_{target}$ set to 10 bits per $40ms$ frame. For global tokens $\bm{G}$, we directly use the continuous embedding  without a quantization as it can be extracted from speech prompt in downstream tasks as described in Section IV, without a need to generate it. As global embedding typically takes a very small portion of the total bitrate in coding, considering its consistency over the long stretches of speech [30], we only count the bits for $\bm{S}$ and $\bm{P}$ in comparison with other acoustic codecs here.  The channel dimensions $C_{g}$, $C_{a}$, $C_{s}$, $C_{p}$ are 256, 768$\times$2, 1024 and 1024, respectively. We use a 12-layer transformer decoder for the generative decoder module, with a hidden dimension of 1024, a feed-forward dimension of 4096, and 8 attention heads. We use adam optimizer with an initial learning rate of 4e-4 for training. The batch size is set to 40 seconds per GPU  and the UniCodec is trained on 16 V100 GPUs.
@@
To evaluate the reconstructed quality, we leverage several metrics that focus on different aspects, including: (i)  Word error rate (WER), which measures the speech intelligibility. we use two public ASR models for English222 https://huggingface.co/facebook/hubert-large-ls960-ft and Spanish  [^3] https://huggingface.co/jonatasgrosman/wav2vec2-large-xlsr-53-spanish, respectively. (ii)  Voicing Decision Error (VDE) [32] and $F_{0}$ Frame Error (FFE) [33], $F_{0}$-related metrics which assess the accuracy of voicing decisions and pitch values. We use them to evaluate the preservation of prosodic information. (iii) SPK-O, cosine similarity of speaker embeddings between the synthesized and the original speech, extracted by a speaker verification model WavLM-TDNN444 https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification. (iv) Emotion-Acc, the top-1 emotion accuracy calculated by an emotion recognition model [^5] https://huggingface.co/ehcalabres/wav2vec2-lg-xlsr-en-speech-emotion-recognition on RAVDESS test set. (v)  The commonly used NISQA [34], UTMOS [35] and VISQOL [36] metrics, which evaluate the naturalness and the overall quality of the synthesized speech. Among them, except NISQA and UTMOS, all others like ASR-WER, VDE, FFE, SPK-O, Emotion-Acc, and VISQOL are metrics with a reference to measure the preservation of various speech information, i.e. the fidelity of reconstructed speech, an indispensable goal for low-bitrate and generative codecs.
@@ V-C Token Representation Capability
TABLE V: Evaluation of zero-shot text-to-speech synthesis on LibriSpeech test-clean with a duration between 4 and 10 seconds. means the results are directly obtained from the paper. means the results are obtained from the official released models. * means the results are obtained from third-party implementations since the official implementation is not publicly available. SPK-D: speaker similarity with compressed prompt. SPK-O: speaker similarity with original prompt. LL-full: the full 60K-hour dataset of Libri-Light. LL-medium: the 6K-hour medium subset of Libri-Light. LT: LibriTTS training data.

### v2

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2024 IEEE
@@ III-B1 Encoder
Instead of training from scratch, we leverage a self-supervised learned semantic encoder for $E_{s}$ and extract residue features through another $E_{p}$ conditioned on intermediate features of $E_{s}$ and $\bm{S}$ (see Fig. 2). Specifically, we initialize $E_{s}$ from a pretrained multilingual HuBERT [5], i.e. mHuBERT Base  [^1] https://dl.fbaipublicfiles.com/hubert/mhubert_base_vp_en_es_fr_it3.pt trained on a 100K subset of the large multilingual corpus VoxPopuli. As the bottom or middle layers of HuBERT capture more acoustic information of speech and the top layers tend to capture more semantic information, we take the $11$th layer output of mHuBERT as our semantic feature $\bm{S}_{se}\ {\in\mathbb{R}^{T\times C_{a}}}$ as that in [19] and the $6$th layer of mHuBERT as our acoustic feature $\bm{A}_{ac}\ {\in\mathbb{R}^{T\times C_{a}}}$ to learn the local pair $(\bm{S},\bm{P})$, $\bm{S}\in\mathbb{R}^{T\times C_{s}},\bm{P}\in\mathbb{R}^{T\times C_{p}}$. Here $C_{a}$, $C_{s}$ and $C_{p}$ are channel dimensions and $T$ is the number of 4040ms frames. As our quantization is performed on a 4040ms frame to produce a more compact token sequence while mHuBERT features are in a 2020ms basis, a reshape to 4040ms by stacking two adjacent frames is conducted before extracting $\bm{S}_{se}$ and $\bm{A}_{ac}$. We should emphasize that as our focus is not on how to learn semantic features, but instead on investigating universal tokens capturing various attributes of speech with compactness and disentanglement that can benefit speech language modeling, here we just use existing mHuBERT in semantic encoder.
@@ III-B3 Decoder
[…]
@@
Acoustic Tokens with TF-Codec We use the pretrained TF-Codec [20]  [^2] https://github.com/microsoft/TF-Codec, a state-of-the-art low-latency neural speech codec, to extract the acoustic token sequence $\bm{A}$. TF-Codec follows a convolutional encoder-VQ-decoder paradigm, operating in the time-frequency domain. For simplicity, we disable the predictive loop in TF-Codec and use the non-predictive pretrained TF-Codec at 6 kbps that achieves high-fidelity speech reconstruction. Unlike RVQ, it uses group vector quantization (GVQ) where the input features stacked at 50 Hz are split into $G_{a}=16$ groups along the channel dimension and each group is quantized by a separate vector quantizer with $1024$ codewords, yielding the acoustic token sequence $\bm{A}=[\bm{a}_{0},\bm{a}_{1},...,\bm{a}_{T_{a}-1}]^{\mathsf{T}}\in\mathbb{R}^{T_{a}\times G_{a}}$. $\bm{a}_{t}=[a_{t}^{0},a_{t}^{1},...,a_{t}^{G_{a}-1}]^{\mathsf{T}}$ denotes the acoustic codes at $t$-th frame with $G_{a}$ groups and each element in $\bm{a}_{t}$ belongs to the set $\{0,1,2,...,1023\}$. $T_{a}$ is the number of acoustic frames.
@@
where $\bm{a}_{-1}$ is the startstart token and $\bm{a}_{T_{a}}$ is the eoseos token. $\bm{a}_{<t}$ denotes all past frames from $\bm{a}_{-1}$ to $\bm{a}_{t-1}$ and we follow this notation in all equations below if not otherwise specified. To achieve this, the $G_{a}$ embeddings corresponding to $G_{a}$ codebooks are concatenated along the channel dimension as the transformer input and $G_{a}$ classification heads are employed to predict the $G_{a}$ acoustic codes in $\bm{a}_{t}$ jointly by projecting the transformer’s final hidden state to $G_{a}$ sets of logits, one for each of the $G_{a}$ codebooks. It is noted that as the number of frames for $\bm{A}$ and $F_{fused}$ are different, i.e. $T_{a}=2T$, a reshape of $F_{fused}$ from $\mathbb{R}^{T\times{C_{a}}}$ to $\mathbb{R}^{T_{a}\times\frac{C_{a}}{2}}$ is conducted before being taken as the condition in the transform decoder.
@@ IV-A Zero-shot Text-to-Speech Synthesis (TTS)
[…]
where $\tilde{x}$ is a speech prompt and texttext is the text input. The $UniCodecLM$ is an autoregressive codec language model conditioned on texttext and $\tilde{x}$, formulated as
@@
[…]
@@ IV-B Speech-to-Speech Translation (S2ST)
$\hskip-2.84526pt\mathcal{L}_{S2ST}\hskip-2.84526pt=\hskip-2.84526pt-\hskip-2.84526pt\ {\sum\limits_{t=0}^{N}\log}p(\bm{c}_{tgt,t}|\bm{c}_{tgt,{<t}},\bm{C}_{src},\bm{L}_{src},\bm{L}_{tgt};\theta_{S2ST}),$ (14)
@@ V-A Speech Resynthesis
[…]
@@
To evaluate the reconstructed quality, we leverage several metrics that focus on different aspects, including: (i) Word error rate (WER), which measures the speech intelligibility. we use two public ASR models for English [^3] https://huggingface.co/facebook/hubert-large-ls960-ft and Spanish  [^4] https://huggingface.co/jonatasgrosman/wav2vec2-large-xlsr-53-spanish, respectively. (ii) Voicing Decision Error (VDE) [32] and $F_{0}$ Frame Error (FFE) [33], $F_{0}$-related metrics which assess the accuracy of voicing decisions and pitch values. We use them to evaluate the preservation of prosodic information. (iii) SPK-O, cosine similarity of speaker embeddings between the synthesized and the original speech, extracted by a speaker verification model WavLM-TDNN [^5] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification. (iv) Emotion-Acc, the top-1 emotion accuracy calculated by an emotion recognition model [^6] https://huggingface.co/ehcalabres/wav2vec2-lg-xlsr-en-speech-emotion-recognition on RAVDESS test set. (v) The commonly used NISQA [34], UTMOS [35] and VISQOL [36] metrics, which evaluate the naturalness and the overall quality of the synthesized speech. Among them, except NISQA and UTMOS, all others like ASR-WER, VDE, FFE, SPK-O, Emotion-Acc, and VISQOL are metrics with a reference to measure the preservation of various speech information, i.e. the fidelity of reconstructed speech, an indispensable goal for low-bitrate and generative codecs.
@@ V-C Token Representation Capability
TABLE V: Evaluation of zero-shot text-to-speech synthesis on LibriSpeech test-clean with a duration between 4 and 10 seconds. means the results are directly obtained from the paper. means the results are obtained from the official released models. * means the results are obtained from third-party implementations since the official implementation is not publicly available. SPK-D: speaker similarity with compressed prompt. SPK-O: speaker similarity with original prompt. $L$-full: the full 60K-hour dataset of Libri-Light. $L$-medium: the 6K-hour medium subset of Libri-Light. LT: LibriTTS training data.